I'm not sure what exactly we are doing but i know its algebra. My question is how do I simplify 3(x+4)-2(x-2)=, plz help me as soon as possible
Answer by tjnw79(57)

3(x+4)-2(x-2)=
frist take the 3 and times what is in the parethesis next to it and do the same with the 2 so it should look like this
3x+12-2x-4
then add you like terms 3x-2x= 1x or (x)
and 12-4=8
so x+8=
if you are solving for x just subtract the 8 on both sides so x=-8